# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ II-C Problem Reformulation
$+\|\mathbf{x}_{i}-\mathbf{y}\|^{2}-2r_{i}\underbrace{\|\mathbf{x}_{i}-\mathbf{y}\|}_{nonsmooth}+r_{i}^{2},\quad\forall i\in\mathcal{N}$ (4)
It is evident that the above loss function contains non-smooth terms. Based on the existing literature[31, 32], we apply the Cauchy-Schwarz inequality to obtain:
@@
$\|\mathbf{x}_{i}-\mathbf{y}\|$ $=\max_{{\mathbf{u}_{i}\in\mathcal{B}}}\mathbf{u}_{i}^{T}(\mathbf{x}_{i}-\mathbf{y}),$ (5b)
where $\mathbf{v}_{i,j},\mathbf{u}_{i}\in\mathbb{R}^{n}$ are auxiliary variables. Then we have
$-\|\mathbf{x}_{i}-\mathbf{x}_{j}\|$ $=\min_{\mathbf{v}_{i,j}\in\mathcal{B}}-\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{x}_{j}),$ (6a)
$-\|\mathbf{x}_{i}-\mathbf{y}\|$ $=\min_{\mathbf{u}_{i}\in\mathcal{B}}-\mathbf{u}_{i}^{T}(\mathbf{x}_{i}-\mathbf{y}).$ (6b)
Substituting (6) into (II-C), the objective function in problem (2a) is rewritten as a minimization problem of a smooth function over a ball constraint set as follows
$\frac{1}{2}\sum_{i\in\mathcal{N}}$ $\Bigg\{\sum_{j\in\mathcal{N}_{i}}\left(\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}+2d_{i,j}\min_{\mathbf{v}_{i,j}\in\mathcal{B}}-\mathbf{v}_{i,j}^{T}\left(\mathbf{x}_{i}-\mathbf{x}_{j}\right)\right)$
$\quad+\left\|\mathbf{x}_{i}-\mathbf{y}\right\|^{2}+2r_{i}\min_{\mathbf{u}_{i}\in\mathcal{B}}-\mathbf{u}_{i}^{T}\left(\mathbf{x}_{i}-\mathbf{y}\right)\Bigg\}$ (7)
This can be conveniently reformulated as the minimization of a smooth function over a simple convex constraint set, as follows:
$\min_{\mathbf{x},\mathbf{y},\mathbf{w}}\sum_{i\in\mathcal{N}}$ $\Bigg\{\sum_{j\in\mathcal{N}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}\left(\mathbf{x}_{i}-\mathbf{x}_{j}\right)\right)$
$\quad\quad+\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}\right\|^{2}-r_{i}\mathbf{u}_{i}^{T}\left(\mathbf{x}_{i}-\mathbf{y}\right)\Bigg\}$ (8)
subject to $\quad\mathbf{w}_{i}\in\mathcal{B}^{N_{i}+1},\forall i\in\mathcal{N},$
where $\mathbf{w}:=\operatorname{vec}(\mathbf{w}_{i},i\in\mathcal{N})$, and for each $i\in\mathcal{N}$, we define $\mathbf{w}_{i}:=\operatorname{vec}(\mathbf{w}_{i,k},\;k\in\{1,\dots,N_{i}+1\})\in\mathbb{R}^{(N_{i}+1)n},\mathbf{v}_{i}:=\operatorname{vec}(\mathbf{v}_{i,j},j\in\mathcal{N}_{i})$. Let the $k$-th element in the set $\mathbf{v}_{i}$ be denoted by $\mathbf{v}_{j_{k}}$. Then, the components of $\mathbf{w}_{i}$ are given by
$\mathbf{w}_{i,k}=\begin{cases}\mathbf{v}_{i,j_{k}},&\text{for }k=1,\dots,N_{i},\\ \mathbf{u}_{i},&\text{for }k=N_{i}+1.\end{cases}$
Let $F_{i}:\mathbb{R}^{(N_{i}+1)n}\times\mathbb{R}^{n}\times\mathbb{R}^{(N_{i}+1)n}\to(-\infty,+\infty]$ be a properly defined lower semicontinuous function, defined as:
$F_{i}(\mathbf{x},\mathbf{y},\mathbf{w}_{i}):=$ $\sum_{j\in\mathcal{N}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}\left(\mathbf{x}_{i}-\mathbf{x}_{j}\right)\right)$
$+\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}\right\|^{2}-r_{i}\mathbf{u}_{i}^{T}\left(\mathbf{x}_{i}-\mathbf{y}\right)$
$+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i}),\quad\forall i\in\mathcal{N}$ (9)
@@
subject to $\quad\mathbf{x}_{k}=\mathbf{a}_{k},\quad\forall k\in\mathcal{A}.$ (10b)
We have the following equality, which allows us to reformulate the problem as in (10):
$\min_{\mathbf{x},\mathbf{y},\mathbf{w}}\sum_{i\in\mathcal{N}}F_{i}(\mathbf{x},\mathbf{y},\mathbf{w}_{i})$
$=$ $\min_{\mathbf{x},\mathbf{y}}\frac{1}{2}\sum_{i\in\mathcal{N}}f_{i}(\mathbf{x},\mathbf{y})-\frac{1}{2}\sum_{i\in\mathcal{N}}(\sum_{j\in\mathcal{N}_{i}}d_{i,j}^{2}+r_{i}^{2}).$ (11)
To solve problem (10) in a distributed manner, we let each node maintain a local estimate of the target position, denoted by $\mathbf{y}_{i}$. A natural way to achieve consensus on the target’s position across the network would be to impose the constraint $\mathbf{y}_{i}=\mathbf{y},\quad\forall i\in\mathcal{N}$, which requires a central node to collect all $\mathbf{y}_{i}$[33]. To avoid this, we instead leverage the connectivity of the sensor network and enforce consistency only between neighboring nodes. Global consensus can then be achieved indirectly by imposing the following constraint:
$\mathbf{y}_{i}=\mathbf{y}_{j},\quad\forall(i,j)\in\mathcal{E}$ (12)
Based on the above derivation, problem (10) can be equivalently rewritten in the form.
$\min_{\mathbf{x},\mathbf{y},\mathbf{w}}\quad\sum_{i\in\mathcal{N}}$ $F_{i}(\mathbf{x},\mathbf{y}_{i},\mathbf{w}_{i}),$ (13a)
$\text{subject to }\quad\mathbf{x}_{k}$ $=\mathbf{a}_{k},\quad\forall k\in\mathcal{A},$ (13b)
$\mathbf{y}_{i}$ $=\mathbf{y}_{j},\quad\forall(i,j)\in\mathcal{E}.$ (13c)
It is observed that the optimization variables $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ ($i\neq j$) are coupled in the loss function (13a), while $\mathbf{y}_{i}$ and $\mathbf{y}_{j}$ ($i\neq j$) are coupled through the consensus constraints (13c),. To enable distributed optimization, we follow standard practices and introduce auxiliary variables that decouple the objective and constraints in problem (10), reformulating it into a form suitable for ADMM. These variables replicate the position variables of neighboring nodes and their target estimates. Specifically, we define: let $\mathbf{p}_{i,j}^{+}\in\mathbb{R}^{n}$ and $\mathbf{q}_{i,j}^{+}\in\mathbb{R}^{n}$ be thevalues of the position $\mathbf{x}_{j}$ of each neighboring node $j\in\mathcal{N}_{i}$ and the estimate $\mathbf{y}_{j}$ of the target position, respectively, i.e.,
$\mathbf{p}_{i,j}^{+}:=\mathbf{x}_{j},\quad j\in\mathcal{N}_{i}.$ (14a)
$\mathbf{q}_{i,j}^{+}:=\mathbf{y}_{j},\quad j\in\mathcal{N}_{i}.$ (14b)
Similarly, let $\mathbf{p}_{i,j}^{-}\in\mathbb{R}^{n}$ and $\mathbf{q}_{i,j}^{-}\in\mathbb{R}^{n}$ be thevalues of the position $\mathbf{x}_{i}$ and the target position estimate $\mathbf{y}_{i}$, respectively, at the neighboring node $j\in\mathcal{N}_{i}$, i.e.,
$\mathbf{p}_{i,j}^{-}:=\mathbf{x}_{i},\quad j\in\mathcal{N}_{i}.$ (14c)
$\mathbf{q}_{i,j}^{-}:=\mathbf{y}_{i},\quad j\in\mathcal{N}_{i}.$ (14d)
@@
By collecting all the variables associated with node $i$, we define a new variable $\mathbf{z}_{i}$ as
$\mathbf{z}_{i}:=\begin{bmatrix}\mathbf{x}_{i}\\ \mathbf{p}_{i}^{-}\\ \mathbf{p}_{i}^{+}\\ \mathbf{y}_{i}\\ \mathbf{q}_{i}^{-}\\ \mathbf{q}_{i}^{+}\end{bmatrix}\in\mathbb{R}^{(4N_{i}+2)n},\quad i\in\mathcal{N}.$ (15)
@@
Here, $\mathbf{p}_{i}^{-}:=\operatorname{vec}(\mathbf{p}_{i,j}^{-},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ and $\mathbf{q}_{i}^{-}:=\operatorname{vec}(\mathbf{q}_{i,j}^{-},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ represent the collections of $N_{i}$values of the position $\mathbf{x}_{i}$ and the target position estimate $\mathbf{y}_{i}$, respectively. Similarly, $\mathbf{p}_{i}^{+}:=\operatorname{vec}(\mathbf{p}_{i,j}^{+},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ and $\mathbf{q}_{i}^{+}:=\operatorname{vec}(\mathbf{q}_{i,j}^{+},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ represent the collections of allvalues of $\mathbf{x}_{j}$ and $\mathbf{y}_{j}$ received from neighboring nodes $j\in\mathcal{N}_{i}$, respectively. Using these notations, Equation (II-C) can be rewritten in a node-separable form as follows:
$\sum_{i\in\mathcal{N}}F_{i}(\mathbf{x},\mathbf{y}_{i},\mathbf{w}_{i})$
@@
$\quad+\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}_{i}\right\|^{2}-r_{i}\mathbf{u}_{i}^{T}\left(\mathbf{x}_{i}-\mathbf{y}_{i}\right)+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i})\bigg]$ (16)
Therefore, based on Equation (II-C), problem (10a) can be reformulated in the following compact form:
$\min_{\mathbf{z},\mathbf{w}}\sum_{i\in\mathcal{N}}\left[\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}+\delta_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}\right)\right]$ (17)
@@
$\mathbf{D}_{i}$ $:=\operatorname{Diag}\left(\operatorname{vec}(d_{i,j},j\in\mathcal{N}_{i}),r_{i}\right)\otimes\mathbf{I}_{n}$
$\in\mathbb{R}^{n(N_{i}+1)\times n(N_{i}+1)}$
represents the measurement matrix of node $i$ and
$\mathbf{H}_{i}$ $:=\begin{bmatrix}\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}$
$\in\mathbb{R}^{n(1+N_{i})\times n(2+4N_{i})}.$ (18)
First, the constraints (12) and (14) can be combined into a compact form:
$\mathbf{A}_{i}\mathbf{z}_{i}=0,$ (19)
where
$\mathbf{A}_{i}$ $:=\begin{bmatrix}\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}$
$\in\mathbb{R}^{n(3N_{i})\times n(2+4N_{i})}.$ (20)
Then, based on (15), the anchor nodes in problem (2) can be represented in the following compact form:
@@
where $\mathcal{X}:=\{\mathbf{z}\mid\mathbf{E}_{i}\mathbf{z}_{i}=\mathbf{a}_{i},\,\forall\,i\in\mathcal{A}\},$ and $\mathbf{E}_{i}:=\begin{bmatrix}1,&\mathbf{0}_{4N_{i}+1}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\times(2+4N_{i})n}.$
Moreover, for a pair of connected sensors $i$ and $j$, we can derive from equations (14c) and (14d) that $\mathbf{p}_{j,i}^{-}=\mathbf{x}_{j}$ and $\mathbf{q}_{j,i}^{-}=\mathbf{y}_{j}$, respectively. Combining these two equations with (14a) and (14b), we obtain the following constraints for all connected sensors:
@@
$\mathcal{Y}_{2}:=\{\mathbf{z}\mid\mathbf{q}_{i,j}^{+}=\mathbf{y}_{j}=\mathbf{q}_{j,i}^{-},\,\forall\,i\in\mathcal{N},\,j\in\mathcal{N}_{i}\}.$

### v2

@@ 2.3 Problem Reformulation
$+{\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}}\big(\|\mathbf{x}_{i}-\mathbf{y}_{k}\|^{2}-2r_{i,{\color[rgb]{0,0,0}k}}\underbrace{\|\mathbf{x}_{i}-\mathbf{y}_{k}\|}_{nonsmooth}+r_{i,{\color[rgb]{0,0,0}k}}^{2}\big)}.$ (4)
It is evident that the above loss function contains non-smooth terms. Based on the existing literature, we apply the Cauchy-Schwarz inequality to obtain:
@@
${\color[rgb]{0,0,0}\|\mathbf{x}_{i}-\mathbf{y}_{k}\|}$ ${\color[rgb]{0,0,0}=\max_{{\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}\in\mathcal{B}}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}(\mathbf{x}_{i}-\mathbf{y}_{k}),}$ (5b)
where $\mathbf{v}_{i,j},{\color[rgb]{0,0,0}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}}\in\mathbb{R}^{n}$ are auxiliary variables. Then the objective function in problem (2a) can be conveniently reformulated as the minimization of a smooth function over a simple convex constraint set, as follows:
$\min_{\mathbf{x},{\color[rgb]{0,0,0}\mathbf{y}},\mathbf{w}}\sum_{i\in\mathcal{N}}$ $\Bigg\{\sum_{j\in\mathcal{N}_{i}}\left(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{x}_{j})\right)$
$+$ ${\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}}\left(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{y}_{k}\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}(\mathbf{x}_{i}-\mathbf{y}_{k})\right)}\Bigg\}$ (6)
subject to $\mathbf{w}_{i}\in\mathcal{B}^{N_{i}+{\color[rgb]{0,0,0}M}},\quad\forall i\in\mathcal{N},$
where $\mathbf{w}:=\operatorname{vec}(\mathbf{w}_{i},i\in\mathcal{N})$. For each $i\in\mathcal{N}$, we define $\mathbf{w}_{i}:=\operatorname{vec}(\mathbf{w}_{i,l},\;l\in\{1,\dots,N_{i}+M\})\in\mathbb{R}^{(N_{i}+M)n}$, where $\mathbf{w}_{i,l}=\mathbf{v}_{i,l}$ for $l=1,\dots,N_{i}$, and $\mathbf{w}_{i,l}=\mathbf{u}_{i,{l-N_{i}}}$ for $l=N_{i}+1,\dots,N_{i}+M$. To ensure a uniform computational structure, dummy variables $\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}$ are included for $k\notin\mathcal{M}_{i}$, which will not affect the objective value.
Let $F_{i}:\mathbb{R}^{(N_{i}+1)n}\times\mathbb{R}^{M_{i}n}\times\mathbb{R}^{(N_{i}+M)n}\to(-\infty,+\infty]$ be a properly defined lower semicontinuous function:
$F_{i}(\mathbf{x},{\color[rgb]{0,0,0}\mathbf{y}},\mathbf{w}_{i})$ (7)
$:=\sum_{j\in\mathcal{N}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}\left(\mathbf{x}_{i}-\mathbf{x}_{j}\right)\right)$
$\quad+{\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}_{k}\right\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}\left(\mathbf{x}_{i}-\mathbf{y}_{k}\right)\right)}$
$\quad+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i}),\quad\forall i\in\mathcal{N}$ (8)
@@
subject to $\quad\mathbf{x}_{l}=\mathbf{a}_{l},\quad\forall l\in\mathcal{A}.$ (9b)
To facilitate a decentralized solution for problem (9), we let each sensor sensor $i\in\mathcal{N}$ maintain its own local estimate of the position of target $\color[rgb]{0,0,0}k\in\mathcal{M}$}, denoted by $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\in\mathbb{R}^{n}$. To ensure that the local estimates across all sensors reach a common consensus on the true target positions without relying on a central controller, we enforce consistency between neighboring sensors in the communication graph:
$\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}=\mathbf{y}_{j,k},\quad\forall(i,j)\in\mathcal{E},\forall{\color[rgb]{0,0,0}k\in\mathcal{M}}.$ (10)
An important feature of this distributed formulation is its ability to handle limited sensing ranges. Specifically, even if target $k$ is not within the line-of-sight (LoS) of sensor $i$ (i.e., $k\notin\mathcal{M}_{i}$), sensor $i$ still maintains and updates the local variable $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}$. Through the iterative collaborative exchange prescribed by (10), target-related information propagates from sensors with direct observations to those without, eventually enabling every sensor in the connected network to estimate the positions of all targets.
For a compact representation, we collect all local estimates at sensor $i$ into a concatenated vector $\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}:=[\mathbf{y}_{i,1}^{T},\dots,\mathbf{y}_{i,M}^{T}]^{T}\in\mathbb{R}^{Mn}$. Furthermore, let $\tilde{\mathbf{y}}:=[\tilde{\mathbf{y}}_{1}^{T},\dots,\tilde{\mathbf{y}}_{N}^{T}]^{T}\in\mathbb{R}^{NMn}$ represent the collection of target estimates across the entire network. Based on these definitions, problem (9) can be equivalently reformulated as the following constrained separable optimization problem:
$\min_{\mathbf{x},{\color[rgb]{0,0,0}\tilde{\mathbf{y}}},\mathbf{w}}\quad$ $\sum_{i\in\mathcal{N}}\Big\{\sum_{j\in\mathcal{N}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}\left(\mathbf{x}_{i}-\mathbf{x}_{j}\right)\right)$
$+{\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\right\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}\left(\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\right)\right)}$
$+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i})\Big\},$ (11a)
subject $\text{ to}\quad\mathbf{x}_{l}=\mathbf{a}_{l},\quad\forall l\in\mathcal{A},$ (11b)
$\quad\quad{\color[rgb]{0,0,0}\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}}={\color[rgb]{0,0,0}\mathbf{y}_{j,k}},\quad\forall(i,j)\in\mathcal{E},{\color[rgb]{0,0,0}\forall{\color[rgb]{0,0,0}k\in\mathcal{M}}},$ (11c)
To decouple the variables for the ADMM framework, we introduce auxiliary variables for both sensor positions and multi-target estimates. Let $\mathbf{p}_{i,j}^{+}\in\mathbb{R}^{n}$ and $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}\in\mathbb{R}^{n}$ be the copies of the position $\mathbf{x}_{j}$ and the target estimate $\mathbf{y}_{j,k}$ maintained by neighbor $j$, respectively. Similarly, let $\mathbf{p}_{i,j}^{-}\in\mathbb{R}^{n}$ and $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}\in\mathbb{R}^{n}$ be the copies of sensor $i$’s own position and target estimates assigned to neighbor $j$:
$\mathbf{p}_{i,j}^{+}:$ $=\mathbf{x}_{j},\quad{\color[rgb]{0,0,0}\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}:=\mathbf{y}_{j,k}},\quad j\in\mathcal{N}_{i},{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}k\in\mathcal{M}}},$ (12a)
$\mathbf{p}_{i,j}^{-}:$ $=\mathbf{x}_{i},\quad{\color[rgb]{0,0,0}\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}:=\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}},\quad j\in\mathcal{N}_{i},{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}k\in\mathcal{M}}}.$ (12b)
@@
By aggregating all variables associated with sensor $i$, we define the high-dimensional local variable $\mathbf{z}_{i}$ as:
$\mathbf{z}_{i}:=\begin{bmatrix}\mathbf{z}^{\text{sen}}_{i}\\ \mathbf{z}^{\text{tar}}_{i}\end{bmatrix}\in\mathbb{R}^{(2N_{i}+1)(1+M)n},\quad i\in\mathcal{N}.$ (13)
Here, $\mathbf{z}^{\text{sen}}_{i}:=[(\mathbf{x}_{i})^{T},(\tilde{\mathbf{p}}_{i}^{-})^{T},(\tilde{\mathbf{p}}_{i}^{+})^{T}]^{T}\in\mathbb{R}^{(2N_{i}+1)n}$ , $\mathbf{z}^{\text{tar}}_{i}:=[(\mathbf{y}_{i,1})^{T},(\tilde{\mathbf{q}}_{i,1}^{-})^{T},(\tilde{\mathbf{q}}_{i,1}^{+})^{T},...,(\mathbf{y}_{i,M})^{T},(\tilde{\mathbf{q}}_{i,M}^{-})^{T},$
@@
$(\tilde{\mathbf{q}}_{i,M}^{+})^{T}]^{T}\in\mathbb{R}^{(2N_{i}+1)Mn}$, where $\mathbf{p}_{i}^{-}:=\operatorname{vec}(\mathbf{p}_{i,j}^{-},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ and $\forall k\in\mathcal{M},\mathbf{q}_{i,{\color[rgb]{0,0,0}k}}^{-}=\operatorname{vec}(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-},j\in\mathcal{N}_{i})$ represent the collections of $N_{i}$values of the position $\mathbf{x}_{i}$ and the target position estimate $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}$, respectively. Similarly, $\mathbf{p}_{i}^{+}:=\operatorname{vec}(\mathbf{p}_{i,j}^{+},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ and $\mathbf{q}_{i,{\color[rgb]{0,0,0}k}}^{+}:=\operatorname{vec}(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+},j\in\mathcal{N}_{i})\in\mathbb{R}^{nN_{i}}$ represent the collections of allvalues of $\mathbf{x}_{j}$ and $\mathbf{y}_{j,k}$ received from neighboring sensors $j\in\mathcal{N}_{i}$, respectively.
Using these notations, Equation (7) can be rewritten in a sensor-separable form as follows:
$\sum_{i\in\mathcal{N}}F_{i}(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{w}_{i})$ (14)
@@
$\quad+{\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}}\left(\frac{1}{2}\left\|\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\right\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}\left(\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\right)\right)}$
$\quad+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i})\bigg].$ (15)
Problem (14) can be rewritten in the following compact form:
$\min_{\mathbf{z},\mathbf{w}}\sum_{i\in\mathcal{N}}\left[\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}\right)\right]$ (16)
where $\mathbf{D}_{i}:=\operatorname{Diag}\left(\operatorname{vec}(d_{i,j},j\in\mathcal{N}_{i}),\operatorname{vec}({\color[rgb]{0,0,0}r_{i,{\color[rgb]{0,0,0}k}},{\color[rgb]{0,0,0}k\in\mathcal{M}}})\right)\otimes\mathbf{I}_{n}$ is the measurement matrix. For algebraic consistency and uniform computation, we define $r_{i,{\color[rgb]{0,0,0}k}}=0$ if target $k\notin\mathcal{M}_{i}$. This convention ensures that the dimensions of $\mathbf{D}_{i}$ and the corresponding mapping matrix $\mathbf{H}_{i}$ remain invariant across all nodes. To this end, The mapping matrix $\mathbf{H}_{i}$ is defined as:
$\mathbf{H}_{i}:=\begin{bmatrix}\mathbf{H}_{i}^{\text{sen}}\\ \mathbf{H}_{i}^{\text{tar}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n(N_{i}+M)\times n(2N_{i}+1)(1+M)}.$ (17)
where $\mathbf{H}_{i}^{\text{sen}}$ captures the relative positions between sensors: $\mathbf{H}_{i}^{\text{sen}}=\begin{bmatrix}\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\dots&\mathbf{0}_{N_{i}}\end{bmatrix}\in\mathbb{R}^{N_{i}\times(2N_{i}+1)(1+M)},$ The target measurement matrix $\mathbf{H}_{i}^{\text{tar}}=\operatorname{vec}(\mathbf{h}_{i,{\color[rgb]{0,0,0}k}},{\color[rgb]{0,0,0}k\in\mathcal{M}})\in\mathbb{R}^{M\times(2N_{i}+1)(1+M)}$ incorporates the row blocks $\mathbf{h}_{i,{\color[rgb]{0,0,0}k}}$ for all $\color[rgb]{0,0,0}k\in\mathcal{M}$}. Specifically, we define the row block $\mathbf{h}_{i,{\color[rgb]{0,0,0}k}}$ such that if target ${\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}$,
$\mathbf{h}_{i,{\color[rgb]{0,0,0}k}}:=\begin{bmatrix}1&0&\dots&0&\underset{\begin{subarray}{c}\uparrow\\ \text{Column for }\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\end{subarray}}{-1}&0&\dots&0\end{bmatrix}$;
otherwise, $\mathbf{h}_{i,{\color[rgb]{0,0,0}k}}=\mathbf{0}_{(2N_{i}+1)(1+M)}^{T}$ for $k\notin\mathcal{M}_{i}$.
The linear constraints (10) and (12) are unified into the compact form
$\mathbf{A}_{i}\mathbf{z}_{i}=\mathbf{0},$ (18)
@@
$\mathbf{A}_{i}:=\operatorname{blockdiag}\left(\mathbf{A}_{i}^{\text{sen}},\mathbf{A}_{i,1}^{\text{tar}},\dots,\mathbf{A}_{i,M}^{\text{tar}}\right)\otimes\mathbf{I}_{n},$ (19)
while the first block handles sensor-related constraints:
$\mathbf{A}_{i}^{\text{sen}}:=\begin{bmatrix}\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\end{bmatrix}\in\mathbb{R}^{N_{i}\times(1+2N_{i})},$
and the subsequent $M$ blocks enforce consistency for each target $\color[rgb]{0,0,0}k\in\mathcal{M}$}:
$\mathbf{A}_{i,{\color[rgb]{0,0,0}k}}^{\text{tar}}:=\begin{bmatrix}\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}\end{bmatrix}\in\mathbb{R}^{2N_{i}\times(1+2N_{i})}.$
Then, based on (13), the anchor sensors in problem (2) can be represented in the following compact form:
@@
where $\mathcal{X}:=\{\mathbf{z}\mid\mathbf{E}_{i}\mathbf{z}_{i}=\mathbf{a}_{i},\,\forall\,i\in\mathcal{A}\},$ and $\mathbf{E}_{i}:=\begin{bmatrix}1,&\mathbf{0}_{{\color[rgb]{0,0,0}(2N_{i}+1)M+2N_{i}}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\times{\color[rgb]{0,0,0}(2N_{i}+1)(1+M)n}}.$
Moreover, for a pair of connected sensors $i$ and $j$, we can derive from equations (12b) that $\mathbf{p}_{j,i}^{-}=\mathbf{x}_{j}$ and $\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-}=\mathbf{y}_{j,k}$, respectively. Combining these two equations with (12a), we obtain the following constraints for all connected sensors:
@@
$\mathcal{Y}_{2}$ $:=\{\mathbf{z}\mid\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}=\mathbf{y}_{j,k}=\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-},\,\forall k\in\mathcal{M},i\in\mathcal{N},\,j\in\mathcal{N}_{i}\}.$